OpenStudy (anonymous):

The main tank has a radius of 70 feet. What is the volume of the quarter-sphered sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (anonymous):

Idon't understand this because its asking for the volume on a 2D object...

OpenStudy (anonymous):

@GingaTheNinja<3 help?

OpenStudy (anonymous):

sorry idk D: but ill recommend ppl c: @ganeshie8 @AccessDenied

OpenStudy (anonymous):

@lollymolly

OpenStudy (anonymous):

XD wow wasn't counting that the sphere is completely circular and the diameter is technically the height XD sorry for the inconvenience .

OpenStudy (anonymous):

Just keeping notes, now.

OpenStudy (anonymous):

area = 15393.8

OpenStudy (anonymous):

15393.8/4 = 3848.45

OpenStudy (anonymous):

whale density = 0.000011142

OpenStudy (anonymous):

3848.45/density?

OpenStudy (anonymous):

3848.45/0.000011142=345400287.202 NO

OpenStudy (anonymous):

3848.45*0.000011142?

OpenStudy (anonymous):

0.0428794299

OpenStudy (anonymous):

1436760 volume

OpenStudy (anonymous):

1436760/4 = (1/4)tank = 359190

OpenStudy (anonymous):

0.000011142

OpenStudy (anonymous):

359190*0.000011142=

OpenStudy (anonymous):

4

OpenStudy (anonymous):

The holding tanks are congruent in size and both are in the shape of a cylinder that has been cut in half vertically. The bottom of the tank is a curved surface. What is the volume of both tanks if the radius of tank #1 is 15 feet and the height of tank #2 is 120 feet? You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (anonymous):

it is similar to the sphere only difference is that it is quarter sized

OpenStudy (anonymous):

i know, I'm just doing this for note taking, its easier to calculate everything i need

OpenStudy (anonymous):

rad = 15 height = 120 diam = ? vol = ? area = ?

OpenStudy (anonymous):

diam = 30

OpenStudy (anonymous):

(3.14*15^2) =

OpenStudy (anonymous):

base = 706.5

OpenStudy (anonymous):

vol = 706.5*120 = 84780

OpenStudy (anonymous):

84780 / 2 = 42390

OpenStudy (anonymous):

holding tanks 1 & 2 vol = 42390

OpenStudy (anonymous):

Next: The stadium is in need of a medical pool to attend to injured or sick whales. The pool will be located adjacent to the holding tanks. Your job is to design a medical pool using a 3-dimensional polyhedral shape you learned about in this module. For safety reasons, the tank can be no more than one-fourth the volume of a holding tank. You will need to submit a 2-dimensional drawing, the dimensions, the volume of your pool, and proof that the pool's volume is no more than one-fourth the volume of a holding tank to your instructor. Step 3: Reflections Step 4: What to Submit

OpenStudy (anonymous):

42390/4=

OpenStudy (anonymous):

max volume allowed for medical pool 10597.5

OpenStudy (anonymous):

volume equation for half sphere = V=((4/3)(3.14)(radius)^3)/2

OpenStudy (anonymous):

>>>((4/3)(3.14)(17)^3)/2<<<